KyledaGreat:

Find the perimeter of the triangle whose vertices are (−1,−6), (14,2), and (8,−6). Write the exact answer. Do not round.

Vocaloid:

sketching the triangle will help visualizing this problem for each side of the triangle, use the distance formula between the endpoints \[d = \sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\] let's use the first two points as an example: let's let (x1,y1) = (−1,−6) and (x2,y2) = (14,2) plugging these into the formula: \[d = \sqrt{(14-(-1))^{2}+(2-(-6))^{2}}\]= sqrt(15^2 + 8^2) = sqrt(289) = 17 repeat with the two other sides using points (−1,−6) and (8,−6) as the second side and points (14,2) and (8,−6) as the third side. add the sides together. like the problem says, do not round.